OpenStudy (anonymous):

Solve equation by factoring. 1) 5v^2=-30-40 Please someone teach me how to do this so I will learn how to do this for the test and thank you!! ^.^

OpenStudy (anonymous):

well for factoring first make sure everything is in the form Ax^2 +By +C =0

OpenStudy (anonymous):

Cant you reduce 70/5?

OpenStudy (anonymous):

Yeah ok iforgot what about that factoring

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Please help I havce a test tomorow

OpenStudy (anonymous):

That Ax^2+B+C=0 thing I forgot how to use it

OpenStudy (anonymous):

is this solving by factoring or completing the square?

OpenStudy (anonymous):

what do you want

OpenStudy (anonymous):

It says solve each equation by factoring.

OpenStudy (anonymous):

Did u jus delete that

OpenStudy (anonymous):

\[5v^2=-30-40\] \[5v^2=-70\] \[v^2=-14\] \[v = -i \sqrt(14)\] \[v = i \sqrt(14)\]

OpenStudy (anonymous):

ya those were bad numbers

OpenStudy (anonymous):

What is the i for? again?

OpenStudy (anonymous):

Did you simplify it some_someone? to get 14? By dividing 5 on both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

But what is the i for though?

OpenStudy (anonymous):

divide by 5. the "i" represents imaginary numbers. or complex numbers.

OpenStudy (anonymous):

Can you explain how that works plz?

OpenStudy (anonymous):

\[i = \sqrt{-1}\] so... \[\sqrt{-x} = i \sqrt{x}\]

OpenStudy (anonymous):

EXAMPLE Ax ^2 + By + C = 0 | | | same form | | | x ^2 + x - 20 = 0

OpenStudy (anonymous):